OpenStudy (anonymous):

se appropriate formulae, but no calculator, to find the exact value sinh(ln(2))of and sinh−1(ln(2)). Show all your steps and express your answers without any powers. do you guys know how to do it???

OpenStudy (anonymous):

shnh= (e^x - e^-x )/ 2 can you do it now?

OpenStudy (anonymous):

sinh(x) = (e^x - e^-x )/ 2 just to be a bit clearer....

OpenStudy (anonymous):

and for the inverse?

OpenStudy (anonymous):

you can find the inverse using what we wrote up there.. anyway : sin-1hx = ln(x+sqrt(x^2 + 1))